OpenStudy (anonymous):

the angles of triangle are x-60,2x-180,xwhat type of triangle is it ?

OpenStudy (anonymous):

The sum of the angles is 180 degrees. So look at \[(x-60) + (2x-180) + x = 180\] solve for x, then work out the other 2 angles. Possible answers would be accute, obtuse, right, equilateral, or isosceles.

OpenStudy (anonymous):

x - 60 + 2x - 180 + x = 180 4x -240 = 180 4x = 420 x = 105 The angles are 105°, (105 -60) = 45º, and (2*105-180) = 30°.